# SPRMamba: Surgical Phase Recognition for Endoscopic Submucosal Dissection with Mamba

### v2

@@ 1 Introduction
    We propose a novel surgical phase recognition framework (SPRMamba), leveraging the Scaled Residual TranMamba (SRTM) module to efficiently model short-term and long-term temporal contexts.
@@ 3.1 Preliminaries
State Space Models (SSMs) are typically considered linear time-invariant systems that map a 1-D function or sequence $x(t)\in\mathbb{R}$to $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{N}$. These systems are commonly represented by linear ordinary differential equations (ODEs) with $A\in\mathbb{R}^{N\times N}$ as the evolution parameter, and $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$ as projection parameters.
@@ 3.2 Overview
The architecture of the proposed SPRMamba is shown in Fig.2. For a video $V\in\mathbb{R}^{L\times C\times H\times W}$ of length ${L}$, we first extract the spatial frame-level feature sequence ${F}\in\mathbb{R}^{L\times D}$ from a fixed ResNet-50 [44], where $D=2048$ is the spatial dimension, and then SPHMamba uses a linear layer to reduce the feature dimension to 64. As in previous works, we repeated each LSTContext block $N$ times, where the dilation factor of the dilation convolution was increased in each layer. In practice, we use $N=10$ and we evaluate the impact of $N$ in Section 4. After the first $N$ layers of the LTContext block, we use an additional linear layer to further reduce the dimensionality $D$ to 32. The dimensionality reduction method reduces the number of parameters from 2.49 million to 1.23 million without degrading the accuracy. We also use an additional Conv1D followed by a softmax layer to generate the frame-level class probabilities ${P}\in\mathbb{R}^{L\times C}$.
We proceed with three additional stages, each consisting of $N$ layers of LSTContext blocks. At the beginning of each stage, we reset the dilation factor of the temporal convolution to 1 and compute the frame-level class probabilities $P\in\mathbb{R}^{T\times C}$ after each stage, resulting in a multi-stage loss. We use cross-entropy loss and the mean squared error smoothing loss introduced by [8] for supervised training. Inspired by [yi2021asformer], we apply the cross-attention to the LSTContext blocks in stages 2 to 4, instead of using feature ${F}$ for queries and keys for window and long-term context samples, but instead using predictive ${P}$. We evaluate the impact of the number of stages in Section 4.
@@ 3.3 Scale Residual TranMamba
Specifically, there are three branches in the TranMamba module. The first branch takes the first quarter portion of the input feature ${F_{in}}$ along the channel dimension as input ${F_{b1}}\in\mathbb{R}^{L\times\frac{C}{4}}$, then expands the dimension to $\lambda$${C}$ by a linear transformation, and then activates it using the SiLU [42] function. The second branch in the TranMamba module has inputs similar to the first branch. It takes the second quarter of the input features ${F_{in}}$ along the channel dimension as input ${F_{b2}}\in\mathbb{R}^{L\times\frac{C}{4}}$. Subsequently, these features are sequentially expanded through the dimensional expansion of the linear layer, Conv1d layer, SSM, and LayerNorm. Afterward, the features extracted from both branches are fused using the Hadamard product, which is designed to capture the long-term temporal context information in this way [30]. The third branch takes the last half of the input features ${F_{in}}$ along the channel dimension as input ${F_{b3}}\in\mathbb{R}^{L\times\frac{C}{2}}$. Subsequently, ${F_{b3}}$ is input into the Attention layer and activated using the GELU function [43], aiming to capture short-term temporal context features. Finally, the short-term temporal features achieved from the third branch will be fused with the previously captured long-term temporal features via a linear mapping to obtain feature ${F_{ls}}$. The entire process takes the following form:
@@ 3.4 Hierarchical Sampling Strateg
Temporal Long-range Sample. For the temporal long-range sample, we sample the input every $G$ to divide $G$ non-overlapping sequences and model the temporal context information of each sequence. In ${G=4}$ shown in Fig.4 b, we model the temporal context information of each of the four downsampled sequences. In general, we model the temporal contexts for $G$ ${F_{b3}}\in\mathbb{R}^{W\times\frac{C}{2}}$ where the ${F_{b1}}$ and ${F_{b2}}$ are the same, i.e., ${F_{b1},F_{b2}}\in\mathbb{R}^{W\times\frac{C}{4}}$. The parameter $G$ provides the flexibility to adjust the sparsity based on the available memory budget, e.g., $G=1$ corresponds to the case where the attention is applied to the entire sequence. In practice, we use $G=64$ and evaluate the impact of $G$ in Section 4.
@@ 4.2 Experiments Design
Evalutation Metrics. For surgical phase recognition, we employed four commonly used evaluation metrics to quantitatively assess model performance, which have been used in previous phase recognition work as well [3, 26, 45, 6]. These metrics include Precision, Recall, Jaccard Index, and Accuracy. Accuracy is defined as the percentage of frames across the entire video correctly predicted to be in their ground truth phase. Given the imbalanced phase presented in the video, Precision, Recall, and Jaccard Index refer to phase-level evaluations, calculated within each phase and then averaged across all phases.
@@ 4.3.1 Result on the ESD385 Dataset
SV-RCNet [26] 75.58$\pm$13.46 81.27$\pm$17.23 70.78$\pm$17.46 60.12$\pm$17.77 41.1G 28.76M
SAHC [48] 86.64$\pm$10.63 86.35$\pm$18.07 83.75$\pm$17.55 75.14$\pm$18.20 9.5G 26.27M
Furube et at. [12] 84.79$\pm$11.58 84.85$\pm$18.75 82.11$\pm$17.52 72.60$\pm$18.25 4.5G 24.69M
AI-Endo [11] 83.38$\pm$12.09 84.74$\pm$17.35 82.17$\pm$17.23 72.09$\pm$17.23 5.7G 24.72M
SPRMamba(Ours) 87.70$\pm$11.03 86.84$\pm$16.07 86.89$\pm$15.64 77.74$\pm$17.60 7.5G 25.42M
@@ 4.3.2 Result on the Cholec80 Dataset
In the Cholec80 dataset, we conducted a performance comparison of our methods with state-of-the-art (SOTA) methods, including Endonet [3], MTRCNet-CL [45], LAST [49], SV-RCNet [26], TeCNO [7], TMRNet [6], Trans-SVNet [9], and SAHC [48]. Please note that we re-implemented TeCNO and Trans-SVNet using the model weights provided in the original manuscript. The results of the other state-of-the-art methods were extracted verbatim from their respective published works.
@@
Method Accuracy(%) Precision(%) Recall(%) Jaccard(%)
Endonet [3]+ 81.70$\pm$4.20 73.70$\pm$16.10 79.60$\pm$7.90 -
MTRCNet-CL [45]+ 89.20$\pm$7.60 86.90$\pm$4.30 88.00$\pm$6.90 -
LAST [49]+ 93.12$\pm$4.71 89.25$\pm$5.49 90.10$\pm$5.45 81.11$\pm$7.62
SV-RCNet [26] 85.30$\pm$7.30 80.70$\pm$7.00 83.50$\pm$7.50 -
TeCNO [7] 89.35$\pm$6.70 83.24$\pm$7.21 81.29$\pm$6.61 70.08$\pm$9.08
TMRNet [6] 90.10$\pm$7.60 90.30$\pm$3.30 89.50$\pm$5.00 79.10$\pm$5.70
Trans-SVNet [9] 90.27$\pm$6.48 85.23$\pm$6.97 82.92$\pm$6.77 72.42$\pm$8.92
SAHC [48] 91.80$\pm$8.10 90.30$\pm$6.40 90.00$\pm$6.40 81.20$\pm$5.50
SPRMamba(Ours) 93.12$\pm$4.58 89.26$\pm$6.69 90.12$\pm$5.61 81.43$\pm$6.90
Our comparison results are presented in Table 3. Our method outperforms the other methods in most of the evaluation metrics, except for the average precision, where it ranks third behind MTRCNet-CL and SAHC. Specifically, for average accuracy, SPRMamba matches the high accuracy of LAST (a multi-task learning method that requires additional information in the form of extra labels) but displays a lower standard deviation of approximately 0.13%. Additionally, for average recall and average Jaccard, SPRMamba achieves the best performance with 90.12% and 81.43%, respectively.
@@ 4.4.2 Hierarchical Sampling Strategy
baseline 86.58$\pm$11.58 86.60$\pm$17.48 84.16$\pm$16.80 75.65$\pm$18.70
STContext SRTM 87.03$\pm$11.73 86.95$\pm$17.57 85.20$\pm$17.32 76.86$\pm$17.82
LTContext SRTM 87.38$\pm$11.15 87.27$\pm$17.20 85.42$\pm$16.43 77.07$\pm$17.89
SPRMamba(Ours) 87.70$\pm$11.03 86.84$\pm$16.07 86.89$\pm$15.64 77.74$\pm$17.60
@@ 5 Disscussion
(1) Long-Term Temporal Modeling with Linear Complexity: By integrating Mamba’s state-space architecture, SPRMamba achieves superior long-range dependency modeling (capturing phases up to 20+ minutes) at 27% lower FLOPs than Transformer-based methods. This linear complexity ($O(L)$ vs. $O(L^{2})$) is critical for real-time ESD analysis, where traditional approaches like MS-TCN [12] and Trans-SVNet [9] fail to balance accuracy and speed (Table 2).

### v3

@@ 1 Introduction
    We propose a novel surgical phase recognition framework (SPRMamba), leveraging the Scaled Residual TranMamba (SRTM) module to efficiently model short-term and long-term temporal contexts. To the best of my knowledge, this is the first research to develop a Mamba-based surgical phase recognition model.
@@ 3.1 Preliminaries
State Space Models (SSMs) are typically considered linear time-invariant systems that map a 1-D function or sequence $x(t)\in\mathbb{R}$ $\to$ $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{N}$. These systems are commonly represented by linear ordinary differential equations (ODEs) with $A\in\mathbb{R}^{N\times N}$ as the evolution parameter, and $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$ as projection parameters.
@@ 3.2 Overview
The architecture of the proposed SPRMamba is shown in Fig.2. For a video $V\in\mathbb{R}^{L\times C\times H\times W}$ of length L{L}, we first extract the spatial frame-level feature sequence ${F}\in\mathbb{R}^{L\times D}$ from a fixed ResNet-50 [40], where $D=2048$ is the spatial dimension, and then SPHMamba uses a linear layer to reduce the feature dimension to 64. As in previous works, we repeated each LSTContext block $N$ times, where the dilation factor of the dilation convolution was increased in each layer. In practice, we use $N=10$ and we evaluate the impact of $N$ in Section 4. After the first $N$ layers of the LTContext block, we use an additional linear layer to further reduce the dimensionality $D$ to 32. The dimensionality reduction method reduces the number of parameters from 2.49 million to 1.23 million without degrading the accuracy. We also use an additional Conv1D followed by a softmax layer to generate the frame-level class probabilities ${P}\in\mathbb{R}^{L\times C}$.
We proceed with three additional stages, each consisting of $N$ layers of LSTContext blocks. At the beginning of each stage, we reset the dilation factor of the temporal convolution to 1 and compute the frame-level class probabilities $P\in\mathbb{R}^{T\times C}$ after each stage, resulting in a multi-stage loss. We use cross-entropy loss and the mean squared error smoothing loss introduced by [8] for supervised training. Inspired by [41], we apply the cross-attention to the LSTContext blocks in stages 2 to 4, instead of using feature F{F} for queries and keys for window and long-term context samples, but instead using predictive P{P}. We evaluate the impact of the number of stages in Section 4.
@@ 3.3 Scale Residual TranMamba
[…]
@@ 3.4 Hierarchical Sampling Strateg
Temporal Long-range Sample. For the temporal long-range sample, we sample the input every $G$ to divide $G$ non-overlapping sequences and model the temporal context information of each sequence. In G=4{G=4} shown in Fig.4 b, we model the temporal context information of each of the four downsampled sequences. In general, we model the temporal contexts for $G$ ${F_{b3}}\in\mathbb{R}^{W\times\frac{C}{2}}$ where the $F_{b1}$} and $F_{b2}$} are the same, i.e., ${F_{b1},F_{b2}}\in\mathbb{R}^{W\times\frac{C}{4}}$. The parameter $G$ provides the flexibility to adjust the sparsity based on the available memory budget, e.g., $G=1$ corresponds to the case where the attention is applied to the entire sequence. In practice, we use $G=64$ and evaluate the impact of $G$ in Section 4.
@@ 4.2 Experiments Design
Evalutation Metrics. For surgical phase recognition, we employed four commonly used evaluation metrics to quantitatively assess model performance. These metrics include Precision, Recall, Jaccard Index, and Accuracy. Accuracy is defined as the percentage of frames across the entire video correctly predicted to be in their ground truth phase. Given the imbalanced phase presented in the video, Precision, Recall, and Jaccard Index refer to phase-level evaluations, calculated within each phase and then averaged across all phases. For the Cholec80 dataset, our evaluation protocol adheres to the benchmark standards established (Relaxed metric) in prior work [23] and aligns with recent approaches [3, 25, 45, 5], considering a result accurate if the temporal discrepancy from the ground truth remains within a 10-second threshold. Furthermore, we consistently utilized the standard metric in other experiments.
@@ 4.3.1 Result on the ESD385 Dataset
[…]
@@ 4.3.2 Result on the Cholec80 Dataset
In the Cholec80 dataset, we conducted a performance comparison of our methods with state-of-the-art (SOTA) methods, including LAST [49], SV-RCNet [25], TeCNO [6], TMRNet [5], Trans-SVNet [7], SAHC [48], SKiT [50], and LoViT [51]. Please note that we re-implemented TeCNO and Trans-SVNet using the model weights provided in the original manuscript. The results of the other state-of-the-art methods were extracted verbatim from their respective published works.
@@
Method Relaxed metric Accuracy(%) Precision(%) Recall(%) Jaccard(%)
[…]
Our comparison results are presented in Table 3. Our method outperforms the other methods in all evaluation metrics. Specifically, for average accuracy, SPRMamba attained an accuracy that exceeds the benchmark set by TeCNO by a margin of 3.77 pp. Moreover, our model demonstrates superior performance over LoViT, the leading model for phase recognition, by a difference of 1.62 pp in accuracy. It also exhibits more consistent performance, as evidenced by a reduced standard deviation in accuracy by roughly 2.12 pp in contrast to TeCNO. Additionally, for average precision, average recall, and average Jaccard, SPRMamba achieves the best performance with 89.26%, 90.12%, and 81.43%, respectively. Beyond standard metrics, SPRMamba also proved to be more effective when evaluated against relaxed metrics.
@@ 4.4.2 Hierarchical Sampling Strategy
[…]
@@ 5 Disscussion
(1) Long-Term Temporal Modeling with Linear Complexity: By integrating Mamba’s state-space architecture, SPRMamba achieves superior long-range dependency modeling (capturing phases up to 20+ minutes) at 27% lower FLOPs than Transformer-based methods (SAHC). This linear complexity ($O(L)$ vs. $O(L^{2})$) is critical for real-time ESD analysis, where traditional approaches like MS-TCN [11] and Trans-SVNet [7] fail to balance accuracy and speed (Table 2).